OpenStudy (anonymous):

Express the complex number in trigonometric form. -2i 2(cos 180° + i sin 180°) 2(cos 90° + i sin 90°) 2(cos 270° + i sin 270°) 2(cos 0° + i sin 0°)

OpenStudy (amorfide):

you can write a complex number as Z let Z=0-2i so Z=-2i you know that the complex number Z can be written Z=r(cos(x)+isin(x))

OpenStudy (amorfide):

r is the magnitude of your complex number

OpenStudy (amorfide):

for example if I have a complex number 3-2i r would be the magnitude so \[r=\sqrt{3^{2}+(-2)^{2}}\]

OpenStudy (amorfide):

you can work out the angle by doing \[\theta= \tan^{-1}(\frac{ b }{ a})\]

OpenStudy (amorfide):

where b is the complex number coefficient and a is the real number

OpenStudy (amorfide):

similarly you could figure this out by using a graph

OpenStudy (michele_laino):

hint: